OpenStudy (anonymous):

We have 0.25 kg of a liquid with a heat capacity of 50 J/K. We also have a heat source at 15 W which increases the temperature of the liquid until the temperature becomes constant. The heat source is then turned of and the temperature of the liquid starts to decrease at a rate of 1.2 K / min. Find the Specific heat capacity.

OpenStudy (anonymous):

What I've tried: E_1 = 0.25 * c * 1.2 = 0.3 * c E_2 = P * t = 15 W * 60 s = 900 J E_1 = E_2 0.3 * c = 900 c = 3000 J / kg K But the answer is 2800 J / kg K